OpenStudy (anonymous):

Calc: Find the force against the vertical end of a tank containing water weighing 62.4 lb/ft^3 as an integral with respect to y The tank end is shaped like the bottom half of a semicircle, with radius 3ft

OpenStudy (anonymous):

You can start with: \[F = \int P\space dA\]

OpenStudy (anonymous):

The area of the base is \(A = \pi R^2/2\), so \(dA = \pi R\space dR\)

OpenStudy (anonymous):

the pressure is \(P = \rho gy\)

OpenStudy (anonymous):

Whoa whoa whoa

OpenStudy (anonymous):

Thats not the equation we were given

OpenStudy (anonymous):

We were given F=int(whA)

OpenStudy (anonymous):

And that pressure equation is different too

OpenStudy (anonymous):

We were given p=wh

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@thomaster @aaronq @e.mccormick

jimthompson5910 (jim_thompson5910):

check out this page http://www.math.ucsb.edu/ugrad/fluid.pdf

OpenStudy (anonymous):

I mean I get parts of it, but I always have the length wrong

OpenStudy (perl):

need help?

OpenStudy (anonymous):

Yes please!!!

OpenStudy (anonymous):

I always think that I have it correct, but I'm usually wrong

OpenStudy (perl):

there are two pressure equations, depending on whether you are given the 'mass' density or the weight density. in this problem you are given the weight density, pound / ft^3 so you dont need to multiply by g

OpenStudy (anonymous):

ok

OpenStudy (perl):

for weight density: $$ P = \rho h$$ for mass density $$P = \rho gh$$

OpenStudy (anonymous):

We weren't even given the second one. The only equations we got were fluid pressure (p)=weight density x height and fluid force=p(weight density)(area)

OpenStudy (anonymous):

And I always get hung up on the area

OpenStudy (anonymous):

Actually I messed up the second equation. Its fluid force=whA

OpenStudy (perl):

Pressure = Force / Area so Force = Area * pressure = A * wh

OpenStudy (anonymous):

We weren't given the first equation either

OpenStudy (anonymous):

I'm confused about how to get A in the formula

OpenStudy (anonymous):

Like for this problem, I did \[\sqrt{9-y^2}dy\] as my area, but apparently its wrong

OpenStudy (perl):

one moment

OpenStudy (anonymous):

ok thanks

OpenStudy (perl):

$$ \Large \Delta F_i = \rho (c - y_i) w(y_i) \Delta y $$

OpenStudy (anonymous):

You're going to have to explain that formula to me

OpenStudy (anonymous):

I moved the axis up 3

OpenStudy (perl):

imagine that the trough end is like a vertical plate that you are submerging in water, (but the vertical plate has the shape of a semicircle

OpenStudy (anonymous):

Down 3

OpenStudy (anonymous):

Sorry!!

jimthompson5910 (jim_thompson5910):

perl, it says "The tank end is shaped like the bottom half of a semicircle" so the rectangular strips should be on the other side of the circle

OpenStudy (anonymous):